# Learning Post-Newtonian Corrections from Numerical Relativity

### v1

@@ Abstract
Accurate modeling of gravitational waveforms from compact binary coalescences remains central to gravitational-wave (GW) astronomy. Post-Newtonian (PN) approximations capture the early inspiral dynamics analytically but break down near merger, while numerical relativity (NR) provides the accurate yet computationally expensive waveforms over limited parameter ranges. We develop a physics-informed neural network (PINN) framework that learns corrections mapping PN dynamics and waveforms to their NR counterparts. As a demonstration of the approach, we use the TaylorT4 PN model as the baseline, and train the network on hybridized NR surrogate data (NRHybSur3dq8) to learn higher-order corrections to the orbital dynamics and waveform modes for nonspinning noneccentric systems. Physically motivated loss terms enforce known limits and symmetries, such as vanishing corrections in the Newtonian limit and suppression of odd-$m$ modes in equal-mass systems, promoting consistent and reliable extrapolation beyond the training region. We simultaneously incorporate corrections that account for the different meaning of mass parameters in PN and NR descriptions. The learned corrections significantly reduce the phase and amplitude error through the inspiral up to about $200M$ before the merger. This approach provides a differentiable and computationally efficient bridge between PN and NR, offering a path toward waveform models that generalize more robustly beyond existing NR datasets.
@@ I Introduction
Figure 1: Real parts of $h_{2,2}$ and $h_{2,1}$ for the highest mass ratio validation case of $q=7.84$ for the final $N$ model for PN to NR correction. The waveform mismatch is reduced from $2\times 10^{-1}$ to $1\times 10^{-6}$. All waveforms are phase-aligned at start. Although the PN waveform agrees well with NR at early inspiral parts, it deviates significantly near merger where both phase and amplitude discrepancies emerge. The NN corrections significantly reduce these discrepancies, allowing the PN waveforms to remain consistent with NR counterparts in the late inspiral regime.
@@
By anchoring the model to known PN dynamics, we effectively embed physical information into the learning process, ensuring that the network’s corrections remain consistent with well-understood analytical behavior. In addition, physics-based loss terms impose constraints on these corrections which regularize the solution and promote physically meaningful extrapolation beyond training region. As a first step forward in this framework, we focus on nonspinning, noneccentric system.
@@ III.1 Overview
For the target NR data, we use the NRHybSur3dq8 Varma:2018mmi surrogate model, which reproduces NR waveforms from the SXS catalog Scheel:2025jct with high fidelity and serves as a computationally efficient proxy for direct NR simulations. To construct the training dataset, we generate a set of nonspinning BBH waveforms using NRHybSur3dq8. Specifically, we sample 100 systems uniformly in mass ratio $q\in[1,8]$, corresponding to symmetric mass ratios $\nu\in[0.109,0.25]$. Each waveform is initialized at a reference $(2,2)$ mode frequency of $f_{2,2}=0.00497[1/M]$, which approximately corresponds to an orbital velocity $v\simeq 0.25$ in the PN expansion. The dataset is again divided into an 80-20 training/validation set using a set of fixed random seeds to ensure reproducibility and consistent generalization.
@@
The neural network typically converges to a waveform mismatch of $10^{-6}$ after $4,000$ epochs. The evolution of the waveform mismatch over training iterations is shown in Fig. 7. The mismatch decrease rapidly during the initial phase of training before gradually stabilizing, indicating that the network efficiently captures the dominant correction structure early on. Moreover, the convergence to a similar level of mismatch across all mass ratio $q$ suggests that the model learns a physically consistent correction rather than overfitting to specific configurations.
@@
Figure 8: Distribution of waveform mismatches for the training and validation sets. Dashed vertical lines represent the mean for each distribution. The close overlap of two distributions indicate that the network achieves comparable accuracy on unseen data and does not overfit to the training set.
The distribution of training and validation errors are shown in Fig. 8. Both distributions peak at similar values and exhibit nearly identical shapes. This close overlap indicates that the network maintains comparable accuracy on unseen data and does not overfit to the training samples. In particular, similarity between the training and validation mismatch distributions indicate that the learned correction generalize well beyond the training examples.
@@ III.2 Loss Function
Figure 9: The history of the total loss and its main components. The steady decline and smooth convergence indicate that the combined loss function yields stable optimization and consistent physical regularization across the training domain.
@@
The evolution of the total loss and its main components during training is shown in Fig. 9. The steady decline and smooth convergence shows that the composite loss yields stable optimization and enforces the desired physical regularization.
@@ III.3 Extrapolation
We use the NRHybSur2dq15 surrogate Yoo:2022erv, which is trained on hybrid PN-NR waveforms up to mass ratio $q=15$, as a proxy for NR to assess extrapolation performance. For reference, we also include the mismatch between NR simulations and NRHybSur2d15. Since the NRHybSur2d15 serves here as the effective “ground truth”, it would not be meaningful for our PINN model to achieve errors below this reference. The shaded region therefore marks the range below which further improvement has no physical significance. While our results are above this region, their close proximity demonstrates that the framework generalizes reasonably well beyond its training region.
@@ III.4 Extending the Model
Building on the model’s demonstrated extrapolation capability, we next investigate how incorporating an additional training data affects its performance. In the original setup, the network was trained on data densely sampled over $q\in[1,8]$. We now introduce a single additional training waveform at $q=15$, well separated from the original training range, and examine how the model’s accuracy changes across intermediate mass ratios.
To isolate the impact of this new data point, we keep the original training dataset unchanged and simply append the $q=15$ waveform generated using NRHybSur2dq15. Starting from the previously trained network parameters, we resume optimization for an additional 1000 epochs using the same loss function and hyperparameters. This warm start allows the network to incorporate new high-$q$ constraints while preserving the structure learned over $q\in[1,8]$. Throughout this extended training, we verify that the model’s performance within the original training region remains stable.
@@ IV Conclusion
Since the model we present here is physics-informed, i.e., it solves PN equations with (small) corrections, one would expect improved extrapolation compared to a more data-driven model like surrogates. We find that the network maintains lower mismatch than the surrogate model (NRHybSur3dq8) outside of the training region, achieving $10^{-4}-10^{-3}$ mismatch levels. Moreover, adding a single additional training point at $q=15$ substantially improves performance across intermediate map ratios, demonstrating that the framework can be efficiently extended with only sparsely sampled data.

### v2

@@ Abstract
Accurate modeling of gravitational waveforms from compact binary coalescences remains central to gravitational-wave (GW) astronomy. Post-Newtonian (PN) approximations capture the early inspiral dynamics analytically but break down near merger, while numerical relativity (NR) provides the accurate yet computationally expensive waveforms over limited parameter ranges. We develop a physics-informed neural network (PINN) framework that learns corrections mapping PN dynamics and waveforms to their NR counterparts. As a demonstration of the approach, we use the TaylorT4 PN model as the baseline, and train the network on a remarkably small dataset of only eight hybridized NR surrogate waveforms (NRHybSur3dq8) to learn higher-order corrections to the orbital dynamics and waveform modes for nonspinning noneccentric systems. Physically motivated loss terms enforce known limits and symmetries, such as vanishing corrections in the Newtonian limit and suppression of odd-$m$ modes in equal-mass systems, promoting consistent and reliable extrapolation beyond the training region. We simultaneously incorporate corrections that account for the different meaning of mass parameters in PN and NR descriptions. The learned corrections significantly reduce the phase and amplitude error through the inspiral up to about $200M$ before the merger. This approach provides a differentiable and computationally efficient bridge between PN and NR, offering a path toward waveform models that generalize more robustly beyond existing NR datasets.
@@ I Introduction
Figure 1: Real parts of $h_{2,2}$ and $h_{2,1}$ for the highest mass ratio validation case of $q=7.93$ for the final $N$ model for PN to NR correction. The waveform mismatch is reduced from $2\times 10^{-1}$ to $1\times 10^{-6}$. All waveforms are phase-aligned at start. Although the PN waveform agrees well with NR at early inspiral parts, it deviates significantly near merger where both phase and amplitude discrepancies emerge. The NN corrections significantly reduce these discrepancies, allowing the PN waveforms to remain consistent with NR counterparts in the late inspiral regime.
@@
By anchoring the model to known PN dynamics, we effectively embed physical information into the learning process, ensuring that the network’s corrections remain consistent with well-understood analytical behavior. In addition, physics-based loss terms impose constraints on these corrections which regularize the solution and promote physically meaningful extrapolation beyond training region. As a first step forward in this framework, we focus on nonspinning, noneccentric system. In this work, we train the model on a small set of eight hybrid NR surrogate waveforms, demonstrating the framework’s potential for data efficient PN to NR calibration.
@@ III.1 Overview
For the target NR data, we use the NRHybSur3dq8 Varma:2018mmi surrogate model, which reproduces NR waveforms from the SXS catalog Scheel:2025jct with high fidelity and serves as a computationally efficient proxy for direct NR simulations. To construct the training dataset, we generate a set of nonspinning BBH waveforms using NRHybSur3dq8. Specifically, we sample ten systems uniformly in symmetric mass ratio $\nu\in[0.109,0.25]$, corresponding to mass ratio $q\in[1,8]$. Each waveform is initialized at a reference $(2,2)$ mode frequency of $f_{2,2}=0.00497[1/M]$, which approximately corresponds to an orbital velocity $v\simeq 0.25$ in the PN expansion.
The dataset is then divided into eight training waveforms and two validation waveforms. [^4] This small set of validation waveforms is a holdover from previous experiments and is not optimized for this problem. Future work could explore more robust validation strategies. These correspond to training mass ratios $q\in\{1.0,1.7,2.2,2.6,4.5,5.4,6.5,8.0\}$ and validation mass ratios $q\in\{3.2,3.8\}$. The training waveforms are used to compute the loss that drives optimization, while the validation waveforms provide an independent validation loss that is used solely for monitoring generalization during training. For the final assessment of model performance, we additionally generate a set of 100 waveforms uniformly sampled in $q\in[1,8]$ and evaluate the trained model on these waveforms. This larger resampled set provides a more robust estimate of the final modeling error and of the model’s generalization quality across the mass ratio range.
@@
As mentioned earlier, we estimate the modeling error by evaluating the network on a set of $100$ surrogate-generated waveforms uniformly sampled in mass ratio $q\in[1,8]$. The evolution of these waveform mismatches over training epochs is shown in Fig. 7. The model quickly reproduces fiducial surrogate waveforms with a mean mismatch of order $10^{-5}$ after roughly $2000$ epochs. Beyond this point, however, the training and validation losses begin to diverge, and the generalization quality degrades. This discrepancy reaches a clear peak at epoch $4000$, as shown in Fig. 8, before beginning to decrease again with a rapid drop in the validation loss. By epoch $8000$, the training and validation losses have converged, and the resulting mismatches settle to a uniform spread at the $10^{-6}$ level across the resampled waveforms.
The $4000$-epoch point coincides with a checkpoint at which we restart the optimization. The warm restart retains the optimized parameters but resets the inverse Hessian approximation and gradient information. To verify that the turnover at $4000$ epoch is not an artifact of restarting, we additionally performed a continuous $8000$-epoch run with no restarts. This run exhibits the same peak and subsequent decline at $4000$ epochs, confirming that the feature is intrinsic to the optimization landscape, rather than the restart. Notably, the checkpoint warm restart run shows better agreement between the training and validation losses by epoch $8000$, suggesting that the restart helps stabilize the late-stage optimization.
@@
Figure 8: The history of the total loss and its main components. The losses decrease rapidly during the first 2000 epochs, after which a divergence between training and validation becomes apparent and peaks near 4000 epochs. A warm restart at this point resets the Hessian approximation, restores alignment between the two losses, and leads to improved generalization in the subsequent optimization.
@@ III.3 Extrapolation
We use the NRHybSur2dq15 surrogate Yoo:2022erv, which is trained on hybrid PN-NR waveforms up to mass ratio $q=15$, as a proxy for NR to assess extrapolation performance. For reference, we also include the mismatch between NR simulations and NRHybSur2d15. Since the NRHybSur2d15 serves here as the effective “ground truth”, it would not be meaningful for our PINN model to achieve errors below this reference. The shaded region therefore marks the range below which further improvement has no physical significance. While our results extend slightly outside of this region, their close proximity demonstrates that the framework generalizes reasonably well beyond its training region.
@@ III.4 Extending the Model
Building on the model’s demonstrated extrapolation capability, we next investigate how incorporating an additional training data affects its performance. In the original setup, the network was trained on eight systems with mass ratios $q\in[1,8]$. We now introduce a single additional training waveform at $q=15$, well separated from the original training range, and examine how the model’s accuracy changes across intermediate mass ratios.
To isolate the impact of this new data point, we keep the original training dataset unchanged and simply append the $q=15$ waveform generated using NRHybSur2dq15. Starting from the previously trained network parameters, we resume optimization for an additional 4000 epochs using the same loss function and hyperparameters. This warm start allows the network to incorporate new high-$q$ constraints while preserving the structure learned over $q\in[1,8]$. Throughout this extended training, we verify that the model’s performance within the original training region remains stable.
@@ IV Conclusion
A notable aspect of this study is that the model achieves these results while being trained on only eight waveforms. This demonstrates the data efficiency made possible by embedding physical structure directly into the model, and suggests that such framework may enable the direct use of NR simulations for training even when only a sparse set of waveforms is available. Since the model we present here is physics-informed, i.e., it solves PN equations with (small) corrections, one would expect improved extrapolation compared to a more data-driven model like surrogates. We find that the network maintains lower mismatch than the surrogate model (NRHybSur3dq8) outside of the training region, achieving $10^{-4}-10^{-3}$ mismatch levels. Moreover, adding a single additional training point at $q=15$ substantially improves performance across intermediate map ratios, demonstrating that the framework can be efficiently extended with only sparsely sampled data.